a jeepney driver drives faster and faster backwards on a straight road from 3 m/s to 12 m/s in a time interval of 25. determine its acceleration?

Answer by Alan3354(69443)

a jeepney driver drives faster and faster backwards on a straight road from 3 m/s to 12 m/s in a time interval of 25. determine its acceleration?
------------
25 what?
a = (12 - 3)/25
a = 9/25 = 0.36 m/sec/unit
backwards might be -0.36, but that depends on the frame of reference